# Halilsoy and Chandrasekhar standing gravitational waves in the linear approximation

### v1

@@ Abstract
Halilsoy and Chandrasekhar cylindrical standing gravitational waves correspond to two different classes of solutions to the vacuum Einstein equations. Both families satisfy the definition of standing gravitational waves proposed by Stephani, but only the latter class fulfills the stricter definition introduced by Chandrasekhar. The aim of this research is to compare both classes of solutions within the linear regime. We discover that the linearized Halilsoy and Chandrasekhar standing waves are gravitational analogues of two different types of electromagnetic polarization standing waves — a phenomenon not previously discussed in the literature for exact solutions to the Einstein equations.
@@ I Introduction
The standing waves play important role in many fields of physics. In general relativity, the nonlinearity of the Einstein equations makes gravitational standing waves an interesting but difficult subject of study. In a cylindrical symmetry, the Einstein equations simplify considerably and the problem of standing gravitational waves may be approached in its full non-linear form.
@@
In this paper, we compare these two classes of standing gravitational waves within the linear regime. This approach allows us to clarify their differences and to find their weak field interpretation as gravitational analogues of electromagnetic polarization standing waves. Furthermore, the linearization sheds light on the nuances concerning the definitions of nodes and antinodes.
@@ II Halilsoy waves
where $\lambda>0$ is a constant, and $J_{0}=J_{0}(\rho/\lambda)$ and $J_{1}=J_{1}(\rho/\lambda)$ are the Bessel functions of the first kind and orders 00 and $1$, respectively.
@@ III Chandrasekhar waves
with the boundary conditions $F(0)=F_{0}$, $F_{,\rho}(0)=0$, where $10<F_{0}<$ 1 is a constant. The constant $\lambda>0$ is a parameter of the solution. No exact non-trivial solutions to the equation above are known so it must be solved numerically. The function $\Psi$ can be determined directly from $F$. The remainning metric functions $\nu$ and $\Omega$ are calculated via quadratures. The Chandrasekhar solutions are regular.
@@ IV.1 Halilsoy waves
After some algebra, the non-zero perturbations of the metric (1) (linear in $\epsilon=A/\sigma^{2}$) can be written as (non-zero components)
@@
and $\rho=\sqrt{x^{2}+y^{2}}$. For $\rho\longrightarrow 0$ non-diagonal components of $h_{\mu\nu}$ vanish.
@@ IV.2 Chandrasekhar waves
We repeat the same procedure for the Chandrasekhar solution. The line element (3) with $F=0$ corresponds to Minkowski spacetime in cylindrical coordinates ($\Psi=1$, $\Omega=0$, and $\nu=const$, where without loss of generality we may assume $\nu=0$). In general, the essential part of the Einstein equations, namely the equation (7), can be solved only numerically. The linearized version takes the form of a Bessel equation of order zero
@@
Following Chandrasekhar [2], we observe that the regularity at the axis $\rho=0$ implies $F^{\prime}(0)=0$, hence we choose $c_{2}=0$. Since $F$ should be small, we set $c_{1}=\epsilon/2$. The linearized Chandrasekhar waves can be derived from $F(\rho)=\frac{\epsilon}{2}J_{0}(\rho/\lambda)+O(\epsilon^{2})$. The non-zero components of the perturbation $h^{\prime}_{\alpha\beta}$ of the metric (3) are
@@
where auxiliary functions $Q_{i}$ are given by the formula (9).
We first confirm that $\Box h^{\prime}_{\mu\nu}=0$ and compute the trace-reversed tensor, then proceed to the TT gauge. We find $h^{\prime}=-2\epsilon Q_{0}$. The trace-reverse tensor $\bar{h}^{\prime}_{\mu\nu}$ has the form (non-zero components)
@@ IV.3 Comparison
The linearized Halilsoy and Chandrasekhar solutions have similar forms in the TT gauge. The essential difference is hidden in the components of the metric related to polarization. The formulas for the Chandrasekhar solution (16), (17), (18) differ from those for the Halilsoy solution (10), (11), (12) solely in their non-diagonal components. The expression $\sinh(\alpha)W$ that appears in the Halilsoy solution is replaced by $Q_{1}$ in the Chandrasekhar solution. The Halilsoy solution for $\alpha=0$ is linearly polarized (the metric function $\omega$ vanishes). The Chandrasekhar solution has always nontrivial polarization term $\Omega$ whose time-dependence is shifted by $\pi/2$ in a phase relative to time-dependence of the term $\omega$ for the Halilsoy solution. The inspection of formulas (12), (18) reveals that far away from the center (the large $x$) both waves become superpositions of plane waves with a supressing factor in amplitudes.
@@ IV.3.1 Test sphere
$\begin{split}\Delta s(\hat{n})=&r\left[1+\frac{\epsilon}{2}[Q_{0}(\cos^{2}\theta-\sin^{2}\theta\sin^{2}\phi)\right.\\ &\left.-K\sin 2\theta\sin\phi-\frac{\lambda}{x_{0}}Q_{1}\sin^{2}\theta\cos 2\phi\right]\;,\end{split}$ (19)
@@
where we used the fact that $\lim_{x_{0}\rightarrow 0^{+}}J_{1}(x_{0})/x_{0}=1/2$. The deformation of a cross-section of the centered test sphere is presented in Fig. 1.
Figure 1: The deformation of a cross-section $y=0$ of a test sphere by the Halilsoy and Chandrasekhar standing gravitational waves. The sphere is centered at the origin and plotted for $t/\lambda\in\{0,\pi/4,\pi/2,3\pi/4,\pi\}$. The thick ring represents the unperturbed sphere at $t/\lambda=\pi/2$. The figure is invariant under rotations in $\phi$ (the horizontal direction). The deformation is exaggerated for clarity ($\epsilon=1/2$).
@@ IV.3.2 Polarization
A cross-section of the test sphere along a plane orthogonal to the direction of propagation of waves ($\phi=\pi/2$) seen as a function of $\rho$ forms a tube. Figs. 2, 3 show how this tube may be deformed by the standing waves (see enclosed animations [9]). For the Halilsoy waves the form of deformations depends strongly on the parameter $\alpha$ as will be explained later in the text.
@@
If $[\dots]=0$ at some $\rho$, then the deformation at that location is independent of $\theta$. In all other cases, the extremal distorisons occur at $\theta_{ex}\in\{0,\pi/2,\pi,-\pi/2\}$, which demonstrates that Halilsoy solution with $\alpha=0$ (i.e., the Einstein-Rosen standing wave) is linearly polarized, as expected.
@@
Figure 2: The deformation of a test tube along $\partial_{\rho}$ direction by the Halilsoy standing gravitational wave. The snapshot is taken at $t/\lambda=7\pi/4$ for $\alpha=2$ and it extends from $\rho/\lambda=2\pi$ to $\rho/\lambda=8\pi$. The black deformed ring at $\rho/\lambda=8\pi$ corresponds to the cross-section of the tube at $t/\lambda=11\pi/4$. The deformation is exaggerated for clarity ($\epsilon=3/2$).
Figure 3: The deformation of a test tube along $\partial_{\rho}$ direction by the Chandrasekhar standing gravitational wave. The snapshot is taken at $t/\lambda=7\pi/4$ and it extends from $\rho/\lambda=2\pi$ to $\rho/\lambda=8\pi$. The black deformed ring at $\rho/\lambda=8\pi$ corresponds to the cross-section of the tube at $t/\lambda=11\pi/4$. The deformation is exaggerated for clarity ($\epsilon=3/2$). The right-handness of the polarization along $\partial_{\rho}$ is clearly visible.
@@
The intermediate values of the parameter $\alpha$ leads to more complicated behaviour. This is a result of a non-trivial mixing of two polarization modes. Figs. 5, 6 show the angles of extremal distortions $\theta_{ex}$ for the Halilsoy standing wave for $\alpha=0.01$ and $\alpha=2$, respectively. The small nonzero parameter $\alpha$ induces a sharp transition, wherein the polarization deviates from the plus mode in a localized region. The growing $\alpha$ extends this transition region which, for large $\alpha$, covers almost the whole space and the cross polarization becomes dominant.
@@
In summary, the Halilsoy solution is plus polarized for $\alpha=0$ (the Einstein-Rosen standing wave). For large $\alpha$ the cross polarization dominates. For intermediate values of $\alpha$ the direction of extremal distortion of the test tube evolves in space and time in a way presented in Figs. 5, 6, 7, 8. The complicated behaviour of the polarization for the intermediate values of $\alpha$ can be explained by the following analysis in the asymptotic region $x\gg\lambda$ near a spatial infinity.
@@
Our analysis shows that the linearized Halilsoy standing wave far from the center corresponds to, using the terminology borrowed from electrodynamics, the “polarization standing wave” [10]. It is a special case of this type of standing wave — the superposition of two orthogonaly linearly polarized waves of equal amplitude propagating in opposite directions. These waves are not identical — the cross polarization term is shifted in a phase by $\pi$.
@@
Let us consider superposition of two linearized plane waves with a circular polarization of the same handedness. We assume that these waves are moving in the opposite directions along $x$-axis. We define these waves in the TT gauge by the following formulas. Non-zero components of the first wave are
@@
The second wave is identical, but it moves in the oposite direction which is equivalent to the substitution $t\rightarrow-t$. The superposition of these two waves forms polarization standing wave. The resulting distortion of the tube is given by the formula
@@ IV.3.3 Nodes and antinodes
The linearized solutions allows us to refine the definitions of “node” and “antinode” in standing gravitational waves. According to a straightforward interpretation, nodes are the timelike hypersurfaces where the distortion of the test tube either vanishes or is minimal, whereas antinodes are the timelike hypersurfaces where the distortion is maximal. Ideally, the nodes would be determined by $t$ and $\theta$ independents zeros of the function $\Delta s(\hat{n})|_{\phi=\pi/2}-r$ and the antinodes would be determined by $t$ and $\theta$ independent extremas of this function. The behaviour the solutions studied in this paper is more complicated and these idealized assumptions, in general, are not satisfied even in the asymptotic region $\rho\gg\lambda$. The minima and extrema of $\Delta s$ depend on $\theta$, so the term “maximal and minimal distortions” needs clarification.
The problem can be approached in several distinct ways. However, the most refined definition we developed for the solutions under investigation is to define the nodes and antinodes in terms of the circumference of a cross-section of the test tube. This definition cannot be applied in the asymptotic region $\rho\gg\lambda$ because, in this limit, the circumference of a test tube become constant. Nevertheless, the analogy with polarization standing waves in electrodynamics gives us a clear understanding of the solutions far from the symmetry center, clarifying in which sense the linearized Halilsoy and Chandrasekhar solutions are standing gravitational waves there.
In vacuum spacetimes, all curvature is encoded entirely in the Weyl tensor. Although the volumes of test spheres remain constant during evolution, neither the area nor circumference of their cross-sections is conserved in general. Consequently, we propose ignoring deformations in other directions and defining nodes and antinodes by the circumference of the test tube’s cross-section in the direction orthogonal to the wave’s direction. Other definitions are possible, e.g. in terms of the area of the cross-section of the tube, but the position of nodes and antinodes defined in this way may fluctuate and be time-independent only on average or in the asymptotic region. Thus, our definition best captures the sense in which the solutions under investigation can be regarded as standing waves.
For both linearized solutions studied in this paper the circumference of the cross-section $C$ of the test tube is given by a remarkably simple formula
$C/r=2\pi+\epsilon\frac{\lambda}{\rho}J_{1}(\rho/\lambda)\cos(t/\lambda)\;.$ (29)
@@
Figure 10: The spatial component of the distortion of the circumference of the test tube $C-2\pi r$ (black) plotted against spatial components of distortions $\Delta s(\theta_{ex})-r$ (gray). The vertical lines correspond to zeros (dashed lines) and extrema (solid lines) of the Bessel function $J_{1}(\rho/\lambda)$. They define nodes (dashed lines) and antinodes (dotted lines) of standing waves for the linearized Halilsoy and Chandrasekhar solutions. The figure is plotted for the Halilsoy solution with $\alpha=0$, but it ilustrates the general behaviour of both solutions.
@@
We plot the spatial amplitudes of the distortions [neglecting the factor $\cos(t/\lambda)$] in Fig. 10. The spatial part of the distortion of the circumference of the tube vanishes roughly at places where the spatial components of the tube distortion have their extrema. Therefore, in our definition, the nodes correspond to places at which the test tube, at first glance, seems to attain its maximal distortion — a fact that is clearly counter-intuitive. This happens because the distortions (e.g. at $\theta\in\{0,\pi\}$ and at $\theta=\pm\pi/2$) have different signs so their contribution to the distortion of the circumference of the tube cancels out. Similarly, the circumferential distortion is maximal at places where the amplitude of the distortions is small, but where they have the same sign. These locations are indentified as antinodes in our definition.

### v2

@@ Abstract
Halilsoy and Chandrasekhar cylindrical standing gravitational waves correspond to two different classes of solutions to the vacuum Einstein equations. Both families satisfy the definition of standing gravitational waves proposed by Stephani, but only the latter class fulfills the stricter definition introduced by Chandrasekhar. The aim of this research is to compare both classes of solutions within the linear regime. We discover that the linearized Halilsoy and Chandrasekhar standing waves are gravitational analogs of two different types of electromagnetic polarization standing waves.
@@ I Introduction
The standing waves play important role in many fields of physics. In general relativity, the nonlinearity of the Einstein equations makes gravitational standing waves an interesting but difficult subject of study. In a cylindrical symmetry, the Einstein equations simplify considerably and the problem of standing gravitational waves may be approached in its full nonlinear form.
@@
In this paper, we compare these two classes of standing gravitational waves within the linear regime. This approach allows us to clarify their differences and to find their weak field interpretation as gravitational analogs of electromagnetic polarization standing waves. Furthermore, the linearization sheds light on the nuances concerning the definitions of nodes and antinodes.
@@ II Halilsoy waves
where $\lambda>0$, $A$, $\alpha$ are constants, and $J_{0}=J_{0}(\rho/\lambda)$ and $J_{1}=J_{1}(\rho/\lambda)$ are the Bessel functions of the first kind and orders 00 and $1$, respectively.
@@ III Chandrasekhar waves
with the boundary conditions $F(0)=F_{0}$, $F_{,\rho}(0)=0$, where $10<F_{0}<$ 1 is a constant. The constant $\lambda>0$ is a parameter of the solution. No exact nontrivial solutions to the equation above are known so it must be solved numerically. The function $\Psi$ can be determined directly from $F$. The remaining metric functions $\nu$ and $\Omega$ are calculated via quadratures. The Chandrasekhar solutions are regular.
@@ IV.1 Halilsoy waves
After some algebra, the nonzero perturbations of the metric (1) (linear in $\epsilon=A/\sigma^{2}$) can be written as (nonzero components)
@@
and $\rho=\sqrt{x^{2}+y^{2}}$. For $\rho\longrightarrow 0$ nondiagonal components of $h_{\mu\nu}$ vanish.
@@ IV.2 Chandrasekhar waves
We repeat the same procedure for the Chandrasekhar solution. The line element (3) with $F=0$ corresponds to Minkowski spacetime in cylindrical coordinates ($\Psi=1$, $\Omega=0$, and $\nu=const$, where without loss of generality we may assume $\nu=0$). In general, the essential part of the Einstein equations, namely the Eq. (7), can be solved only numerically. The linearized version takes the form of a Bessel equation of order zero
@@
Following Chandrasekhar [2], we observe that the regularity at the axis $\rho=0$ implies $F^{\prime}(0)=0$, hence we choose $c_{2}=0$. Since $F$ should be small, we set $c_{1}=\epsilon/2$. The linearized Chandrasekhar waves can be derived from $F(\rho)=\frac{\epsilon}{2}J_{0}(\rho/\lambda)+O(\epsilon^{2})$. The nonzero components of the perturbation $h^{\prime}_{\alpha\beta}$ of the metric (3) are
@@
where auxiliary functions $Q_{i}$ are given by the Eq. (9).
We first confirm that $\Box h^{\prime}_{\mu\nu}=0$ and compute the trace-reversed tensor, then proceed to the TT gauge. We find $h^{\prime}=-2\epsilon Q_{0}$. The trace-reverse tensor $\bar{h}^{\prime}_{\mu\nu}$ has the form (nonzero components)
@@ IV.3 Comparison
The linearized Halilsoy and Chandrasekhar solutions have similar forms in the TT gauge. The essential difference is hidden in the components of the metric related to polarization. The formulas for the Chandrasekhar solution (16), (17), (18) differ from those for the Halilsoy solution (10), (11), (12) solely in their nondiagonal components. The expression $\sinh(\alpha)W$ that appears in the Halilsoy solution is replaced by $Q_{1}$ in the Chandrasekhar solution. The Halilsoy solution for $\alpha=0$ is linearly polarized (the metric function $\omega$ vanishes). The Chandrasekhar solution has always nontrivial polarization term $\Omega$ whose time-dependence is shifted by $\pi/2$ in a phase relative to time-dependence of the term $\omega$ for the Halilsoy solution. The inspection of formulas (12), (18) reveals that far away from the center (the large $x$) both waves become superpositions of plane waves with a supressing factor in amplitudes.
@@ IV.3.1 Test sphere
$\begin{split}\Delta s(\hat{n})=&r\left[1+\frac{\epsilon}{2}\big{(}Q_{0}(\cos^{2}\theta-\sin^{2}\theta\sin^{2}\phi)\right.\\ &\left.-K\sin 2\theta\sin\phi-\frac{\lambda}{x_{0}}Q_{1}\sin^{2}\theta\cos 2\phi\big{)}\right]\;,\end{split}$ (19)
@@
where we used the fact that $\lim_{x_{0}\rightarrow 0^{+}}J_{1}(x_{0})/x_{0}=1/2$. The deformation of a cross section of the centered test sphere is presented in Fig. 1.
Figure 1: The deformation of a cross section $y=0$ of a test sphere by the Halilsoy and Chandrasekhar standing gravitational waves. The sphere is centered at the origin and plotted for $t/\lambda\in\{0,\pi/4,\pi/2,3\pi/4,\pi\}$. The thick ring represents the unperturbed sphere at $t/\lambda=\pi/2$. The figure is invariant under rotations in $\phi$ (the horizontal direction). The deformation is exaggerated for clarity ($\epsilon=1/2$).
@@ IV.3.2 Polarization
A cross section of the test sphere along a plane orthogonal to the direction of propagation of waves ($\phi=\pi/2$) seen as a function of $\rho$ forms a tube. Figures 2, 3 show how this tube may be deformed by the standing waves (see enclosed animations [9]). For the Halilsoy waves the form of deformations depends strongly on the parameter $\alpha$ as will be explained later in the text.
@@
If $[\dots]=0$ at some $\rho$, then the deformation at that location is independent of $\theta$. In all other cases, the extremal distorsions occur at $\theta_{ex}\in\{0,\pi/2,\pi,-\pi/2\}$, which demonstrates that Halilsoy solution with $\alpha=0$ (i.e., the Einstein-Rosen standing wave) is linearly polarized, as expected.
@@
Figure 2: The deformation of a test tube along $\partial_{\rho}$ direction by the Halilsoy standing gravitational wave. The snapshot is taken at $t/\lambda=7\pi/4$ for $\alpha=2$ and it extends from $\rho/\lambda=2\pi$ to $\rho/\lambda=8\pi$. The black deformed ring at $\rho/\lambda=8\pi$ corresponds to the cross section of the tube at $t/\lambda=11\pi/4$. The deformation is exaggerated for clarity ($\epsilon=3/2$).
Figure 3: The deformation of a test tube along $\partial_{\rho}$ direction by the Chandrasekhar standing gravitational wave. The snapshot is taken at $t/\lambda=7\pi/4$ and it extends from $\rho/\lambda=2\pi$ to $\rho/\lambda=8\pi$. The black deformed ring at $\rho/\lambda=8\pi$ corresponds to the cross section of the tube at $t/\lambda=11\pi/4$. The deformation is exaggerated for clarity ($\epsilon=3/2$). The right-handedness of the polarization along $\partial_{\rho}$ is clearly visible.
@@
The intermediate values of the parameter $\alpha$ leads to more complicated behavior. This is a result of a nontrivial mixing of two polarization modes. Figures 5, 6 show the angles of extremal distortions $\theta_{ex}$ for the Halilsoy standing wave for $\alpha=0.01$ and $\alpha=2$, respectively. The small nonzero parameter $\alpha$ induces a sharp transition, wherein the polarization deviates from the plus mode in a localized region. The growing $\alpha$ extends this transition region which, for large $\alpha$, covers almost the whole space and the cross polarization becomes dominant.
@@
In summary, the Halilsoy solution is plus polarized for $\alpha=0$ (the Einstein-Rosen standing wave). For large $\alpha$ the cross polarization dominates. For intermediate values of $\alpha$ the direction of extremal distortion of the test tube evolves in space and time in a way presented in Figs. 5, 6, 7, 8. The complicated behavior of the polarization for the intermediate values of $\alpha$ can be explained by the following analysis in the asymptotic region $x\gg\lambda$ near a spatial infinity.
@@
Our analysis shows that the linearized Halilsoy standing wave far from the center corresponds to, using the terminology borrowed from electrodynamics, the “polarization standing wave” [10]. It is a special case of this type of standing wave—the superposition of two orthogonally linearly polarized waves of equal amplitude propagating in opposite directions. These waves are not identical—the cross polarization term is shifted in a phase by $\pi$.
@@
Let us consider superposition of two linearized plane waves with a circular polarization of the same handedness. We assume that these waves are moving in the opposite directions along $x$-axis. We define these waves in the TT gauge by the following formulas. Nonzero components of the first wave are
@@
The second wave is identical, but it moves in the opposite direction which is equivalent to the substitution $t\rightarrow-t$. The superposition of these two waves forms polarization standing wave. The resulting distortion of the tube is given by the formula
@@ IV.3.3 Nodes and antinodes
The linearized solutions allows us to refine the definitions of “node” and “antinode” in standing gravitational waves. According to a straightforward interpretation, nodes are the timelike hypersurfaces where the distortion of the test tube either vanishes or is minimal, whereas antinodes are the timelike hypersurfaces where the distortion is maximal. Ideally, the nodes would be determined by $t$ and $\theta$ independent zeros of the function $\Delta s(\hat{n})|_{\phi=\pi/2}-r$ and the antinodes would be determined by $t$ and $\theta$ independent extremas of this function. The behavior the solutions studied in this paper is more complicated and these idealized assumptions, in general, are not satisfied even in the asymptotic region $\rho\gg\lambda$. The minima and extrema of $\Delta s$ depend on $\theta$, so the term “maximal and minimal distortions” needs clarification.
The problem can be approached in several distinct ways. However, the most refined definition we developed for the solutions under investigation is to define the nodes and antinodes in terms of the circumference of a cross section of the test tube. This definition cannot be applied in the asymptotic region $\rho\gg\lambda$ because, in this limit, the circumference of a test tube become constant. Nevertheless, the analogy with polarization standing waves in electrodynamics gives us a clear understanding of the solutions far from the symmetry center, clarifying in which sense the linearized Halilsoy and Chandrasekhar solutions are standing gravitational waves there.
In vacuum spacetimes, all curvature is encoded entirely in the Weyl tensor. Although the volumes of test spheres remain constant during evolution, neither the area nor circumference of their cross sections is conserved in general. Consequently, we propose ignoring deformations in other directions and defining nodes and antinodes by the circumference of the test tube’s cross section in the direction orthogonal to the wave’s direction. Other definitions are possible, e.g. in terms of the area of the cross section of the tube, but the position of nodes and antinodes defined in this way may fluctuate and be time-independent only on average or in the asymptotic region. Thus, our definition best captures the sense in which the solutions under investigation can be regarded as standing waves.
For both linearized solutions studied in this paper the circumference of the cross section $C$ of the test tube is given by a remarkably simple formula
$C/r=2\pi+\epsilon\frac{\pi}{2}\frac{\lambda}{\rho}J_{1}(\rho/\lambda)\cos(t/\lambda)\;.$ (29)
@@
Figure 10: The spatial component of the distortion of the circumference of the test tube $C/r-2\pi$ (black) plotted against spatial components of distortions $\Delta s(\theta_{ex})/r-1$ (gray). The vertical lines correspond to zeros (dashed lines) and extrema (solid lines) of the Bessel function $J_{1}(\rho/\lambda)$. They define nodes (dashed lines) and antinodes (dotted lines) of standing waves for the linearized Halilsoy and Chandrasekhar solutions. The figure is plotted for the Halilsoy solution with $\alpha=0$, but it illustrates the general behavior of both solutions. The deformation is exaggerated for clarity ($\epsilon=1/2$).
@@
We plot the spatial amplitudes of the distortions [neglecting the factor $\cos(t/\lambda)$] in Fig. 10. The spatial part of the distortion of the circumference of the tube vanishes roughly at places where the spatial components of the tube distortion have their extrema. Therefore, in our definition, the nodes correspond to places at which the test tube, at first glance, seems to attain its maximal distortion—a fact that is clearly counterintuitive. This happens because the distortions (e.g. at $\theta\in\{0,\pi\}$ and at $\theta=\pm\pi/2$) have different signs so their contribution to the distortion of the circumference of the tube cancels out. Similarly, the circumferential distortion is maximal at places where the amplitude of the distortions is small, but where they have the same sign. These locations are identified as antinodes in our definition.